f(x)=2x^2-x+7 ; g(x)=x-1
find (fog)(X)
I don't even know what this means

Answer by stanbon(48535)

f(x)=2x^2-x+7 ; g(x)=x-1
find (fog)(X)
----------------
fog(x) = f[g(x)]
=f[x-1]
= 2(x-1)^2-(x-1)+7
= 2(x^2-2x+1) -x+8
= 2x^2-5x+10
=================
Cheers,
Stan H.